please provide a complete solution and explanation , thank you

Ka. 3. If 4-point signal x[nl-(4,3,2,1), then xI ((-3))4] a)I b) 2 c) 3 Ed)at e) none above
Answer #1

nin) When we shift lineally, the umaining Jampled ave cansiden as eos Whne asin cioculㆂ shift the some signal Will λ1peat atthe signal can be repeated circularly with the given values

From the above figure we can write x(-3)=3

Another method:

X(n)={4,3,2,1}\\ X(n-1)={1,4, 3,2}\\ X(n-2)={2,1,4,3}\\ X(n-3)={3,2,1,4} Put n=0 in x(n-3) implies x(-3)=3